# Large Language Models Struggle in Token-Level Clinical Named Entity Recognition

### v1

@@ Introduction
Recent advancements in computational linguistics have led to the emergence of Large Language Models (LLMs). These models have shown remarkable capabilities in various domains, including the clinical field 12; 13; 14. Their strengths in in-context learning, such as retrieval-augmented generation (RAG), present new opportunities for tackling the challenges of NER in clinical texts. Despite the promise of LLMs, most existing research in this area has either focused on document-level NER (i.e., aggregation NER), which involves identifying entities at a broader level in entire documents or paragraphs without predicting the exact span or location of these entities, or on adapting ChatGPT to this task. For example, Zhou et al. introduce UniversalNER, a distillation approach using mission-focused instruction-tuning to create efficient models that excel in NER. They distill ChatGPT into a more compact LLaMA-based 15 model UniversalNER and demonstrate superior document-level NER accuracy across diverse domains, including biomedicine 16. Similarly, in their exploration of ChatGPT-3.5-turbo, Shry et al. focus on the extraction of rare diseases and their associated phenotypes via prompt engineering, also limiting their analysis to document-level output, which overlooks the precise span and location of these entities 17. In contrast, the potential of LLMs, especially local open-source LLMs, in token-level NER (i.e., span-based NER) in clinical settings, particularly for rare diseases, remains relatively unexplored.
@@
This study aims to explore the effectiveness of LLMs in token-level NER for rare diseases within clinical texts. By focusing on this specific application, we seek to understand the limitations and capabilities of LLMs in handling the challenges of token-level entity recognition in a domain where data is scarce and highly specialized. Essentially, we explore the performance of various state-of-the-art local open-source LLMs, including LLaMA-2 18, Meditron 19, Llama2-MedTuned 20, and UniversalNER 16, on the task of NER of rare diseases and their phenotypes. Additionally, we assess the performance of prominent models such as ChatGPT-3.5 and ChatGPT-4. Our investigation extends to examining their capabilities in in-context learning, evaluating performance enhancements through few-shot learning, retrieval-augmented generation (RAG), and fine-tuning. The experimental results reveal that without proper fine-tuning, local LLMs generally struggle with token-level NER, despite prior training on clinical texts. We observe that while few-shot learning can effectively improve the performance of most LLMs, the influence of RAG on the same task is relatively minimal. Notably, our findings indicate that a medically-adapted LLaMA-2-7b model, specifically Llama2-MedTuned 20, can outperform ChatGPT-4 on this task. This is particularly noteworthy as it achieves these results without having been trained specifically on the rare disease data used in the experiments, and this highlights the potential of fine-tuning local LLMs for specialized applications in clinical NER and other clinical NLP tasks.
@@
Recently, researchers have released various LLMs dedicated to the medical field. LLaMA 15 and LLaMA-2 18 are one of the first and most popular open-source LLMs, laying the foundation for extensive research and applications. Meditron 19 is a suite of open-source medical LLMs pre-trained on the GAP-Replay corpus, including clinical guidelines, research papers, and general domain data, and surpass the performance of LLaMA-2 18, ChatGPT-3.5, and Flan-PaLM 21 on multiple medical reasoning benchmarks. BioMistral 22 is another significant contribution, built upon the Mistral model 23 and further enhanced with training on PubMed Central, showing superior performance on various medical question-answering tasks.
In parallel, considerable efforts have been devoted to adapting these LLMs to the task of NER through prompting or fine-tuning. For example, Zhou et al. instruction-fine-tune the LLaMA model 15 using ChatGPT-generated synthetic data for NER from broad-coverage unlabeled web text 16. Their UniversalNER model shows promising NER performance across multiple domains. However, they only generate document-level outputs, i.e., a list of extracted entities in the given text as shown in Figure 2(a), without considering their exact span information, which limits their impact and usage in practical scenarios. Hu et al. explore the token-level NER capabilities of ChatGPT-3.5 and ChatGPT-4 on two clinical NER tasks by manually crafting specific prompts 24. However, they limit their exploration to ChatGPT models and do not investigate local and open-source LLMs. Another study that aligns closely with our research is by Shry et al., who focus on extracting rare diseases and their phenotypes at a document-level by prompting ChatGPT-3.5 17.
@@ Methods
In this study, we aim to understand the capabilities of various general and medical LLMs in token-level NER, which is underexplored in clinical settings. Similar to Shry et al.’s work 17, we choose rare diseases as our case study. Our focus is particularly on extracting information about rare diseases and their associated phenotypes, driven by two primary reasons: 1) Data related to rare diseases is scarce and highly specialized, presenting an ideal scenario to leverage the strengths of LLMs; and 2) Patients with rare diseases are a relatively understudied group, necessitating increased research and advocacy efforts, as highlighted by Nguyen et al. 25 To this end, our investigation spans across the overall performance of both local open-source LLMs and ChatGPT models on this task. More specifically, we initially explore the zero-shot performance of these LLMs using manually-designed prompts. Secondly, to enhance the models’ performance for this specific task, we employ in-context learning strategies, i.e., few-shot learning and retrieval-augmented generation (RAG). These approaches provide rich information to the prompts to facilitate a deeper understanding of the task by the LLMs. Finally, we instruction-fine-tune Llama2-MedTuned 20 on the rare disease dataset to assess the performance of LLMs in a fully supervised way. Our intention is to explore how well LLMs, when fine-tuned with direct supervision, can adapt to and accurately identify entities in complex rare disease data.
@@
  • LLaMA-2-7b 18 is one of the first and most popular local open-source pre-trained LLMs released. Though it is not dedicated to the clinical domain, the model is often considered a critical benchmark in recent related studies due to its impact and generalizability. For consistency, we utilize the 7-billion-parameter version of all local LLMs in this study.
  • Meditron-7b 19 is an adapted version of LLaMA-2 to the medical domain through continued pre-training on a comprehensively curated medical corpus, including selected PubMed papers and abstracts, a new dataset of internationally-recognized medical guidelines, and general domain data from RedPajama-v1 28. It outperforms LLaMA-2-7B and PMC-LLaMA 29 on multiple medical reasoning tasks.
  • UniversalNER-7b 16 is an instruction-fine-tuned version of LLaMA-7b 15 which is specified in named entity recognition. Essentially, the authors prompt ChatGPT to generate instruction-tuning data for NER from web text and conduct instruction-tuning on LLaMA. UniversalNER shows state-of-the-art document-level NER performance on multiple benchmarks across diverse domains, including biomedicine.
  • Llama2-MedTuned-7b 20 is a medically-adapted version of the LLaMA-2-7b 18 model. Essentially, the authors specifically create a medical instruction-based dataset consisting of approximately $200,000$ samples covering clinical NLP tasks including token-level named entity recognition, relation extraction, document classification, question answering, natural language inference, and conduct instruction-tuning on LLaMA. Llama2-MedTuned demonstrates on-par performance with domain-specific models like BioClinicalBERT 8 on several benchmarks.
@@
In this study, we consider both in-context learning and instruction-fine-tuning to further adapt the LLMs to the token-level NER task. For in-context learning, we employ few-shot learning and retrieval-augmented generation (RAG), the prompt templates of which are shown in Figure 2. Essentially, we randomly select $5$ samples from the training set as the demonstration to the LLMs. To ensure consistency, we fix the $5$ samples for all few-shot experiments in the study. For RAG implementation, we utilize the LlamaIndex framework  [^3] https://github.com/run-llama/llama_index, incorporating NORD rare disease articles as the knowledge base. Specifically, these articles are segmented into chunks ($chunk\_size=512$), embedded using the bge-large-en-v1.5 model 30, and stored in a vector database. In the retrieval stage, we convert a query using the embedding model and retrieve the top-$k$ ($similarity\_top\_k=2$) most relevant text chunks based on similarity with the query embedding. The retrieved chunks are then integrated into the LLM prompts for enriched context.
For instruction-fine-tuning, we follow previous work 20 and convert the RareDis dataset to the Stanford Alpaca 31 format. This format is structured for instruction-following, comprising three parts: an instruction, an input, and an output, as shown in Figure 2. We then fine-tune the Llama2-MedTuned-7b model on the transformed RareDis dataset. We use LoRA adapters 32 for the fine-tuning. It is worth noting that while instruction-fine-tuning (or instruction-tuning) somewhat contradicts our initial intent due to its reliance on sufficient data, it remains crucial for our investigation into LLMs’ performance capabilities and limitations in comparison to smaller BERT-like models in data-rich scenarios.
@@
We utilize Precision (P), Recall (R), and F1-score (F1) as our evaluation metrics for the token-level NER task. For models designed for document-level NER, such as UniversalNER-7b 16, we adapt their outputs to token-level by exact string matching. Initially, we assess a range of LLMs to investigate their overall performance. This is followed by a more focused evaluation of those LLMs that demonstrate proficiency in token-level NER. Subsequently, we conduct a detailed analysis of the aforementioned learning strategies, examining their influence on the performance of the LLMs. Additionally, we perform an error analysis to gain insights into the common errors across different models and learning strategies.
@@ Results
Besides zero-shot prompting and in-context learning, we also investigate the impact of instruction-fine-tuning on these LLMs in this task. As shown in Table 2 and Table 3, despite the performance gain via few-shot learning, the overall performance of prompting LLMs is far from satisfactory. In this experiment, we aim to investigate LLMs’ performance capabilities and limitations compared to fine-tuned BERT-like models under data-rich conditions, which is the common solution to NER in current practice. The results are shown in Figure 3. Essentially, Figure 3 visualizes the F1 scores of various models, including BioClinicalBERT fine-tune, Llama2-MedTuned under different learning strategies (zero-shot, few-shot, fine-tune), and ChatGPT-4 in zero-shot and few-shot scenarios. Specifically, the performance of Llama2-MedTuned models demonstrates significant effectiveness, not only outperforming ChatGPT-4 in most scenarios but also closely rivaling BioClinicalBERT when fine-tuned. This highlights the promising potential of local open-source LLMs in this specialized NER task.
@@
There are also a few options to adapt these LLMs to the token-level NER task. A typical solution is continuous pre-training, or instruction-fine-tuning, just as Llama2-MedTuned and what we do using the rare disease data. Basically, the idea is to cast the NER problem as text generation, so the LLM can have a better understanding of the task and thus process it correctly in the same manner as it was pre-trained. For instance, Yang et al. integrate Human Phenotype Ontology (HPO) labels into clinical notes and fine-tune GPT-based models (e.g., GPT-J) for phenotype recognition 33. They also conduct a comprehensive comparative analysis of encoder-based and decoder-based models for this task. Another solution is to leverage LLMs to generate synthetic data and use the data to train smaller specified models. In addition to the data, one can also alter the model architecture of LLMs. For example, Li et al. propose to fine-tune LLaMA-2 models in a label-supervised way, by removing the causal mask from the decoders to enable bidirectional self-attention which is essential for token classification tasks like NER 34. However, in our preliminary experiment, the method does not work as expected and the performance is worse than instruction-fine-tuning on the rare disease dataset. We leave it for future study.

### v2

@@ Introduction
Recent advancements in computational linguistics have led to the emergence of Large Language Models (LLMs). These models have shown remarkable capabilities in various domains, including the clinical field 12. Their strengths in in-context learning, such as retrieval-augmented generation (RAG), present new opportunities for tackling the challenges of NER in clinical texts. Despite the promise of LLMs, most existing research in this area has either focused on document-level NER (i.e., aggregation NER), which involves identifying entities at a broader level in entire documents or paragraphs without predicting the exact span or location of these entities, or on adapting ChatGPT to this task. For example, Zhou et al. introduce UniversalNER, a distillation approach using mission-focused instruction-tuning to create efficient models that excel in NER. They distill ChatGPT into a more compact LLaMA-based 13 model UniversalNER and demonstrate superior document-level NER accuracy across diverse domains, including biomedicine 14. Similarly, in their exploration of ChatGPT-3.5-turbo, Shry et al. focus on the extraction of rare diseases and their associated phenotypes via prompt engineering, also limiting their analysis to document-level output, which overlooks the precise span and location of these entities 15. In contrast, the potential of LLMs, especially local open-source LLMs, in token-level NER (i.e., span-based NER) in clinical settings, particularly for rare diseases, remains relatively unexplored.
@@
This study aims to explore the effectiveness of LLMs in token-level NER for rare diseases within clinical texts. By focusing on this specific application, we seek to understand the limitations and capabilities of LLMs in handling the challenges of token-level entity recognition in a domain where data is scarce and highly specialized. Essentially, we explore the performance of various state-of-the-art local open-source LLMs, including LLaMA-2 16, Meditron 17, Llama2-MedTuned 18, and UniversalNER 14, on the task of NER of rare diseases and their phenotypes. Additionally, we assess the performance of prominent models such as ChatGPT-3.5 and ChatGPT-4. Our investigation extends to examining their capabilities in in-context learning, evaluating performance enhancements through few-shot learning, retrieval-augmented generation (RAG), and fine-tuning. The experimental results reveal that without proper fine-tuning, local LLMs generally struggle with token-level NER, despite prior training on clinical texts. We observe that while few-shot learning can effectively improve the performance of most LLMs, the influence of RAG on the same task is relatively minimal. Notably, our findings indicate that a medically-adapted LLaMA-2-7b model, specifically Llama2-MedTuned 18, can outperform ChatGPT-4 on this task. This is particularly noteworthy as it achieves these results without having been trained specifically on the rare disease data used in the experiments, and this highlights the potential of fine-tuning local LLMs for specialized applications in clinical NER and other clinical NLP tasks.
@@
Recently, researchers have released various LLMs dedicated to the medical field. LLaMA 13 and LLaMA-2 16 are one of the first and most popular open-source LLMs, laying the foundation for extensive research and applications. Meditron 17 is a suite of open-source medical LLMs pre-trained on the GAP-Replay corpus, including clinical guidelines, research papers, and general domain data, and surpass the performance of LLaMA-2 16, ChatGPT-3.5, and Flan-PaLM 19 on multiple medical reasoning benchmarks.
In parallel, considerable efforts have been devoted to adapting these LLMs to the task of NER through prompting or fine-tuning 13; 14; 15; 20; 21; 22; 23; 24. For example, Zhou et al. instruction-fine-tune the LLaMA model 13 using ChatGPT-generated synthetic data for NER from broad-coverage unlabeled web text 14. Their UniversalNER model shows promising NER performance across multiple domains. However, they only generate document-level outputs, i.e., a list of extracted entities in the given text as shown in Figure 2(a), without considering their exact span information, which limits their impact and usage in practical scenarios. Hu et al. explore the token-level NER capabilities of ChatGPT-3.5 and ChatGPT-4 on two clinical NER tasks by manually crafting specific prompts 20. However, they limit their exploration to ChatGPT models and do not investigate local and open-source LLMs. Another study that aligns closely with our research is by Shry et al., who focus on extracting rare diseases and their phenotypes at a document-level by prompting ChatGPT-3.5 15.
@@ Methods
In this study, we aim to understand the capabilities of various general and medical LLMs in token-level NER, which is underexplored in clinical settings. Similar to Shry et al.’s work 15, we choose rare diseases as our case study. Our focus is particularly on extracting information about rare diseases and their associated phenotypes, driven by two primary reasons: 1) Data related to rare diseases is scarce and highly specialized, presenting an ideal scenario to leverage the strengths of LLMs; and 2) Patients with rare diseases are a relatively understudied group, necessitating increased research and advocacy efforts, as highlighted by Nguyen et al. 25 To this end, our investigation spans across the overall performance of both local open-source LLMs and ChatGPT models on this task. More specifically, we initially explore the zero-shot performance of these LLMs using manually-designed prompts. Secondly, to enhance the models’ performance for this specific task, we employ in-context learning strategies, i.e., few-shot learning and retrieval-augmented generation (RAG). These approaches provide rich information to the prompts to facilitate a deeper understanding of the task by the LLMs. Finally, we instruction-fine-tune Llama2-MedTuned 18 on the rare disease dataset to assess the performance of LLMs in a fully supervised way. Our intention is to explore how well LLMs, when fine-tuned with direct supervision, can adapt to and accurately identify entities in complex rare disease data.
@@
  • LLaMA-2-7b 16 is one of the first and most popular local open-source pre-trained LLMs released. Though it is not dedicated to the clinical domain, the model is often considered a critical benchmark in recent related studies due to its impact and generalizability. For consistency, we utilize the 7-billion-parameter version of all local LLMs in this study.
  • Meditron-7b 17 is an adapted version of LLaMA-2 to the medical domain through continued pre-training on a comprehensively curated medical corpus, including selected PubMed papers and abstracts, a new dataset of internationally-recognized medical guidelines, and general domain data from RedPajama-v1 28. It outperforms LLaMA-2-7B on multiple medical reasoning tasks.
  • UniversalNER-7b 14 is an instruction-fine-tuned version of LLaMA-7b 13 which is specified in named entity recognition. Essentially, the authors prompt ChatGPT to generate instruction-tuning data for NER from web text and conduct instruction-tuning on LLaMA. UniversalNER shows state-of-the-art document-level NER performance on multiple benchmarks across diverse domains, including biomedicine.
  • Llama2-MedTuned-7b 18 is a medically-adapted version of the LLaMA-2-7b 16 model. Essentially, the authors specifically create a medical instruction-based dataset consisting of approximately $200,000$ samples covering clinical NLP tasks including token-level named entity recognition, relation extraction, document classification, question answering, natural language inference, and conduct instruction-tuning on LLaMA. Llama2-MedTuned demonstrates on-par performance with domain-specific models like BioClinicalBERT 8 on several benchmarks.
@@
In this study, we consider both in-context learning and instruction-fine-tuning to further adapt the LLMs to the token-level NER task. For in-context learning, we employ few-shot learning and retrieval-augmented generation (RAG), the prompt templates of which are shown in Figure 2. Essentially, we randomly select $5$ samples from the training set as the demonstration to the LLMs. To ensure consistency, we fix the $5$ samples for all few-shot experiments in the study. For RAG implementation, we utilize the LlamaIndex framework, incorporating NORD rare disease articles as the knowledge base. Specifically, these articles are segmented into chunks ($chunk\_size=512$), embedded using the bge-large-en-v1.5 model, and stored in a vector database. In the retrieval stage, we convert a query using the embedding model and retrieve the top-$k$ ($similarity\_top\_k=2$) most relevant text chunks based on similarity with the query embedding. The retrieved chunks are then integrated into the LLM prompts for enriched context.
For instruction-fine-tuning, we follow previous work 18 and convert the RareDis dataset to the Stanford Alpaca 29 format. This format is structured for instruction-following, comprising three parts: an instruction, an input, and an output, as shown in Figure 2. We then fine-tune the Llama2-MedTuned-7b model on the transformed RareDis dataset. We use LoRA adapters 30 for the fine-tuning. It is worth noting that while instruction-fine-tuning (or instruction-tuning) somewhat contradicts our initial intent due to its reliance on sufficient data, it remains crucial for our investigation into LLMs’ performance capabilities and limitations in comparison to smaller BERT-like models in data-rich scenarios. Further implementation details are available in our code repository [^3] https://github.com/qiuhaolu/tner.
@@
We utilize Precision (P), Recall (R), and F1-score (F1) as our evaluation metrics for the token-level NER task. For models designed for document-level NER, such as UniversalNER-7b 14, we adapt their outputs to token-level by exact string matching. Initially, we assess a range of LLMs to investigate their overall performance. This is followed by a more focused evaluation of those LLMs that demonstrate proficiency in token-level NER. Subsequently, we conduct a detailed analysis of the aforementioned learning strategies, examining their influence on the performance of the LLMs. Additionally, we perform an error analysis to gain insights into the common errors across different models and learning strategies.
@@ Results
Besides zero-shot prompting and in-context learning, we also investigate the impact of instruction-fine-tuning on these LLMs in this task. As shown in Table 2 and Table 3, despite the performance gain via few-shot learning, the overall performance of prompting LLMs is far from satisfactory. In this experiment, we aim to investigate LLMs’ performance capabilities and limitations compared to fine-tuned BERT-like models under data-rich conditions, which is the common solution to NER in current practice. The results are shown in Figure 3. Essentially, Figure 3 visualizes the F1 scores of various models, including BioClinicalBERT fine-tune, Llama2-MedTuned under different learning strategies (zero-shot, few-shot, fine-tune), and ChatGPT-4 in zero-shot and few-shot scenarios. Specifically, the performance of Llama2-MedTuned models demonstrates significant effectiveness, not only outperforming ChatGPT-4 in most scenarios but also closely rivaling BioClinicalBERT when fine-tuned. This highlights the potential of open-source LLMs in this task.
@@
There are also a few options to adapt these LLMs to the token-level NER task. A typical solution is continuous pre-training, or instruction-fine-tuning, just as Llama2-MedTuned and what we do using the rare disease data. Basically, the idea is to cast the NER problem as text generation, so the LLM can have a better understanding of the task and thus process it correctly in the same manner as it was pre-trained. For instance, Yang et al. integrate Human Phenotype Ontology (HPO) labels into clinical notes and fine-tune GPT-based models (e.g., GPT-J) for phenotype recognition 31. They also conduct a comprehensive comparative analysis of encoder-based and decoder-based models for this task. Another solution is to leverage LLMs to generate synthetic data and use the data to train smaller specified models. In addition to the data, one can also alter the model architecture of LLMs. For example, Li et al. propose to fine-tune LLaMA-2 models in a label-supervised way, by removing the causal mask from the decoders to enable bidirectional self-attention which is essential for token classification tasks like NER 32. However, in our preliminary experiment, the method does not work as expected and the performance is worse than instruction-fine-tuning on the rare disease dataset. We leave it for future study.